OpenStudy (anonymous):

What is the point-slope form of the equation passing through (5, -1) with a slope of 6? How do I find it? Medal and fan!

OpenStudy (aum):

Point-Slope form is:\[ y - y_1 = m(x-x_1) \]m is the slope and \((x_1,y_1)\) is the given point.

OpenStudy (anonymous):

y--1=6(x-5) ?

OpenStudy (anonymous):

Yeah, it's right, Thanks! :D